Helllo,
I need help with this word problem:
Mrs. Lambert invested a portion of $15,000 at 10% annual interest and the balance at 6% annual interest. If her investments earned $1260 for the year, how much did she invest in each account?
Thank you!!

Answer by checkley75(3666)

.10X+.06(15000-X)=1260
.10X+900-.06X=1260
.04X=1260-900
.04X=360
X=360/.04
X=9000 INVESTED @ 10%.
15000-9000=6000 INVESTED @ 6%.
PROOF
.10*9000+.06*6000=1260
900+360=1260
1260=1260